### v1

# Distributed Hypothesis Testing Under A Covertness Constraint

Abstract
We study distributed hypothesis testing under a covertness constraint in the non-alert situation, which requires that under the null-hypothesis an external warden be unable to detect whether communication between the sensor and the decision center is taking place. We characterize the achievable Stein exponent of this setup when the channel from the sensor to the decision center is a partially-connected discrete memoryless channel (DMC), i.e., when certain output symbols can only be induced by some of the inputs. The Stein-exponent in this case, does not depend on the specific transition law of the DMC and equals Shalaby and Papamarcou’s exponent without a warden but where the sensor can send $k$ noise-free bits to the decision center, for $k$ a function that is sublinear in the observation length $n$. For fully-connected DMCs, we propose an achievable Stein-exponent and show that it can improve over the local exponent at the decision center. All our coding schemes do not require that the sensor and decision center share a common secret key, as commonly assumed in covert communication. Moreover, in our schemes the divergence covertness constraint vanishes (almost) exponentially fast in the obervation length $n$, again, an atypical behaviour for covert communication.
I Introduction
In distributed hypothesis testing, two distant terminals, a remote sensor and a decision center, observe correlated sources, see Figure 1 (without the external warden). The sources’ underlying joint distribution depends on one of two hypotheses $\mathcal{H}=0$ or $\mathcal{H}=1$ and the goal of the decision center is to guess this underlying hypothesis based on its own observations and communication from the sensor. The performance of the decision center is characterized by two error probabilities [1]: the Type-I error probability of declaring $\hat{\mathcal{H}}=1$ while $\mathcal{H}=0$, denoted $\alpha_{n}$, and the Type-II error probability of declaring $\hat{\mathcal{H}}=0$ while $\mathcal{H}=1$, denoted $\beta_{n}$.
An important line of work in information-theory aims at characterizing the largest exponential decay rate of the Type-II error probability $\beta_{n}$ under the constraint that the Type-I error probability $\alpha_{n}$ lies below a given threshold $\epsilon$ [2, 3, 4, 5, 6, 7, 8]. This largest exponent is commonly referred to as Stein-exponent. It was mostly studied in a centralized setup or when communication from the sensor to the decision center is over a noiseless link. Even in this latter case, the Stein-exponent remains generally an open problem, with the notable exception of certain source distributions [3, 9] or when the number of transmitted bits scales only sublinearly in the observation length [10].
In a recent line of work [11, 12], we extended the results in [10] to a setup where the sensor communicates to the decision center using a discrete memoryless channels (DMC) a number of times that is sublinear in the observation length. Such a scenario arises naturally in practical scenarios where sensors nowadays are often highly energy-limited. Our results [11] showed a dichotomy of the Stein-exponent with respect to the connectivity of the DMC. Whenever the DMC is fully-connected, i.e., each input symbol can induce each output symbol, then the exponent is no better as the local exponent at the decision center. Communication from the sensor thus becomes useless in terms of Stein-exponent. In contrast, when the DMC is only partially-connected, i.e., some outputs can only be induced by some inputs, then the same exponent is achievable as over a noiseless link [10]. We showed that the same result also remains valid when the sensor can use the DMC $n$ (the observation length) times, but a stringent sublinear cost-constraint is imposed.
In following-up work [12], we relaxed the cost-constraint to an expected sublinear cost constraint, and showed a similar dichotomy of the Stein-exponent with respect to the connectivity of the DMC. For partially-connected networks, the Stein-exponent still coincides with the noise-less link exponent in [10]. For fully-connected networks, the relaxed expected cost constraint however allows to improve over the local Stein-exponent at the decision center.
Figure 1: Distributed hypothesis testing with an external warden.
In this work, the sensor again communicates to the decision center over $n$ uses of a DMC. However, under $\mathcal{H}=0$, communication needs to be such that an external warden cannot detect the mere fact that communication is happening. Depending on the practical situation, it might be important that communication be covert under both hypothesie, or only in the case where there is no alert situation, i.e., under $\mathcal{H}=0$. In this paper, we consider the latter scenario. Implicitly, such a covertness constraint imposes that under $\mathcal{H}=0$, the sensor employs most of the time a dedicated zero-symbol that also represents the absence of communication. It can be shown that the covertness constraint is weaker than imposing that the channel can only be used a sublinear number of times but stronger than the expected cost constraint. However, as is common for covert communication, here we allow the sensor and the decision center to share a uniform secret key that is unknown to the warden, see Figure 1.
In this article we show that for some partially-connected DMCs, the Stein-exponent is the same as for the noiseless link with sublinear communication setup in [10], irrespective of the size of the shared secret key. Moreover, it can be achieved without using the secret key at all, and the classical divergence covertness measure [13, 14, 15] can be made to vanish arbitrarily close to exponentially fast in the observation length $n$. These results are in contrast to classical covert communication where the shared secret key helps in improving performance and the covertness constraint vanishes more slowly [14].
For fully-connected DMCs we show that the Stein-exponent can improve over the local exponent at the decision center. Our coding and testing scheme again does not use the shared secret key and achieves divergence covertness metrics that vanish exponentially fast in the blocklength $n$. All our results show that the Stein-exponent does not depend on the imposed Type-I error thresholds $\epsilon$.
Our setup has previously been considered in [16], where it was shown that the noiseless link exponent provides an upper bound on the Stein-exponent under a covertness constraint. The proof however only holds when the key length grows logarithmically in $n$. In this paper we extend this upper bound for arbitrary key lengths. Moreover, the work in [16] did not establish any achievability result improving over the local exponent.
It is worth mentioning also the works in [17, 18, 19, 20] where privacy constraints are imposed on a distributed hypothesis testing setup. It is required that a warden that eavesdrops on the perfedct noiseless link from the source to the receiver be unable to tell certain parts of the the sensor’s source observations. For example, [19] proved the Stein-exponent under such a privacy constraint for source distributions termed “testing against conditional independence” and when the Type-I error probability is required to vanish. Recently, [20] extended this result to arbitrary constraints $\epsilon$ on the Type-I error probability and showed that the Stein exponent depends on $\epsilon$.
Notation: We mostly follow standard notation. Random variables are denoted by uppercase letters (e.g., $X$), while their realizations are denoted by lowercase letters (e.g., $x$). We abbreviate $(x_{1},\ldots,x_{n})$ by $x^{n}$ and $(x_{t+1},\ldots,x_{n})$ by $x_{t+1}^{n}$. The Hamming weight and Hamming distance are denoted by $w_{\textnormal{H}}(\cdot)$ and $d_{\textnormal{H}}(\cdot,\cdot)$, respectively. We further abbreviate independent and identically distributed as i.i.d. and probability mass function as pmf.
We denote by $\pi_{x^{n}y^{n}}$ the joint type of the sequences $(x^{n},y^{n})$, defined as
$\pi_{x^{n}y^{n}}(a,b)\triangleq\frac{n_{x^{n},y^{n}}(a,b)}{n},$ (1)
where $n_{x^{n},y^{n}}(a,b)$ is the number of occurrences of the pair $(a,b)$ in $(x^{n},y^{n})$. We use $\mathcal{T}_{\mu}^{(n)}(P_{XY})$ to denote the jointly strongly typical set as defined in [21, Definition 2.9].
II Problem Setup
Consider the hypothesis testing problem in Figure 1, where the sensor observes a sequence $U^{n}$, and a secret key $S$ and communicates to a decision center, which also knows the secret key $S$ in addition to its local observations $V^{n}$ . Under the null hypothesis
$\mathcal{H}_{0}:(U^{n},V^{n})\ \textnormal{i.i.d.}\ P_{UV}$ (2)
whereas under the alternative hypothesis
$\mathcal{H}_{1}:(U^{n},V^{n})\ \textnormal{i.i.d.}\ Q_{UV}.$ (3)
Similarly to [10] we assume that $P_{UV}(u,v)=0$ whenever $Q_{UV}(u,v)=0$.
The sensor sends an input sequence
$X^{n}=f^{(n)}(U^{n},S)\in\mathcal{X}^{n}$ (4)
over the channel, where $f^{(n)}$ is a chosen encoding function on appropriate domains. The decision center observes the corresponding outputs $Y^{n}$ of the discrete memoryless channel (DMC) $\Gamma_{YZ|X}$ and the warden the outputs $Z^{n}$.
It is assumed that the input alphabet $\mathcal{X}$ contains the 0 symbol. In fact, the covertness constraint imposes that under $\mathcal{H}=0$ the pmf of the warden’s output distribution
$P_{Z^{n}|\mathcal{H}=0}=\frac{1}{\mathcal{K}}\sum_{s\in\mathcal{K}}\sum_{u^{n}\in\mathcal{U}^{n}}P_{U}^{\otimes n}(u^{n})\Gamma_{Z|X}^{\otimes n}(\cdot|f^{(n)}(u^{n},s)),$ (5)
be close to the output distribution $\Gamma_{Z|X=0}^{\otimes n}$ induced by the all-zero input sequence. Specifically, covertness is measured by the Kullback-Leibler divergence:
$d_{n}:=D(P_{Z^{n}|\mathcal{H}=0}||\Gamma_{Z|X=0}^{\otimes n})$ (6)
and is required to stay below a given threshold $\epsilon_{n}$ for sufficiently large blocklengths $n$. Following standard convention in covert communication, we require that the DMC $\Gamma_{YZ|X}$ satisfies the following conditions:
[…]
$\textnormal{Supp}\left(\Gamma_{Z|X}(\cdot|x)\right)$ $\subseteq$ $\textnormal{Supp}\left(\Gamma_{Z|X}(\cdot|0)\right),\quad\forall x\in\mathcal{X},$ (7b)
$\textnormal{Supp}\left(\Gamma_{Y|X}(\cdot|x)\right)$ $\subseteq$ $\textnormal{Supp}\left(\Gamma_{Y|X}(\cdot|0)\right),\quad\forall x\in\mathcal{X}$ (7c)
[…]
Based on the received sequence $Y^{n}$ and its observations $V^{n}$, the decision center produces a guess of the hypothesis:
$\hat{\mathcal{H}}=g^{(n)}(V^{n},Y^{n})\in\{0,1\}.$ (8)
The goal is to design encoding and decision functions such that the Type-I error probability
$\alpha_{n}\triangleq\Pr\left[\hat{\mathcal{H}}=1|\mathcal{H}=0\right]$ (9)
stays below given threshold $\epsilon\geq 0$ and the Type-II error probability
$\beta_{n}\triangleq\Pr\left[\hat{\mathcal{H}}=0|\mathcal{H}=1\right]$ (10)
decays to 0 with largest possible exponential decay, and under hypothesis $\mathcal{H}_{0}:$
$\lim_{n\rightarrow\infty}d_{n}=0.$ (11)
Definition 1.
Given $\epsilon\in[0,1)$, a miss-detection error exponent $\theta>0$ is called $\epsilon$-achievable under a covertness constraint if there exists a sequence of encoding and decision functions $\{(f^{(n)},g^{(n)})\}_{n=1}^{\infty}$ satisfying (11) and
[…]
The supremum over all $\epsilon$-achievable miss-detection error exponents $\theta$ is denoted $\theta_{\textnormal{covert}}^{\star}(\epsilon)$ and called covert Stein-exponent.
The achievable error exponent depends on the transition law $\Gamma_{YZ|X}$ and on the source pmfs $P_{UV}$ and $Q_{UV}$.
III Results
Define the following three exponents:
$E_{1}$ $\triangleq$ $\min_{\begin{subarray}{c}\pi_{UV}\colon\\ \pi_{U}=P_{U}\\ \pi_{V}=P_{V}\end{subarray}}D(\pi_{UV}\|Q_{UV})$ (13a)
$E_{2}$ $\triangleq$ $\min_{\begin{subarray}{c}\pi_{UV}:\ \\ \pi_{V}=P_{V},\;\pi_{U}\notin\bar{\mathcal{P}}_{U}(x_{1})\end{subarray}}D\!\big(\pi_{UV}\,\|\,Q_{UV}\big)$ (13b)
$E_{3}$ $\triangleq$ $\max_{\begin{subarray}{c}x_{1}\in\mathcal{X}\setminus\{0\}\end{subarray}}\Big[\min_{\begin{subarray}{c}\pi_{UV}:\ \\ \pi_{V}=P_{V},\;\pi_{U}\in\bar{\mathcal{P}}_{U}\end{subarray}(x_{1})}D\!\big(\pi_{UV}\,\|\,Q_{UV}\big)$ (13c)
$\hskip 85.35826pt+\ D(\Gamma_{Y|X=0}\|\Gamma_{Y|X=x_{1}})\Big],$
[…]
$\bar{\mathcal{P}}_{U}(x_{1}):=\Big\{\pi_{U}\in\mathcal{P}(\mathcal{U}):\begin{aligned} &D\!\left(\pi_{U}\,\middle\|\,P_{U}\right)\geq\max_{\pi_{z}\in\mathcal{P}(\mathcal{Z})}\Big[D\!\left(\pi_{z}\,\middle\|\,\Gamma_{Z|X}(\cdot|0)\right)-\tfrac{3}{2}D\!\left(\pi_{z}\,\middle\|\,\Gamma_{Z|X}(\cdot|x_{1})\right)\Big]\end{aligned}\Big\}.$ (14)
Theorem 1.
Fix $\epsilon\in[0,1)$.
  1. 1.
    If the DMC is such that there exists an input $\hat{x}\in\mathcal{X}\setminus\{0\}$ and an output $y^{*}\in\mathcal{Y}$ satisfying the two conditions:
    $\Gamma_{Y|X}(y^{*}|0)$ $>$ $0$ (15a)
[…]
    then the miss-detection error probability is given by:
    $\theta_{\textnormal{covert}}^{*}(\epsilon)$ $=$ $E_{1}.$ (16)
  2. 2.
    Otherwise
    $\theta_{\textnormal{covert}}^{*}(\epsilon)$ $\geq$ $\min\big\{E_{2},E_{3}\big\}.$ (17)
Our achievability results do not require that the sensor and decision center use the shared secret key $S$.
The covertness constraint $d_{n}$ can be made to vanish exponentially fast in $n$ for the result in (17) and arbitrary close to exponentially fast in $n$ for the achievability result in (16).
Remark 1.
Exponent $E_{1}$ is also the Stein-exponent in a distributed setup without warden where the sensor can send a sublinear (in $n$) number of noise-free bits to the decision center [10]. For partially-connected DMCs satisfying (15), the covertness constraint (11) thus has the same impact as limiting communication to a sublinear (in $n$) number of noise-free bits.
Remark 2 (Improvement over the Local Test).
Consider the pmf
$T_{U}(u):=\sum_{v}P_{V}(v)\,Q_{U|V}(u\mid v),\quad u\in\mathcal{U}.$ (18)
We observe that if $T_{U}\notin\bar{\mathcal{P}}_{U}(x_{1})$ for all $x_{1}$, then $E_{2}=D(P_{V}\|Q_{V})$, which coincides with the local exponent at the decision center without communication from the sensor.
In particular, $E_{2}=D(P_{V}\|Q_{V})$ when $T_{U}=P_{U}$ (as is the case for testing against independence where $Q_{UV}=P_{U}P_{V}$). For $T_{U}=P_{U}$ also $E_{1}=D(P_{V}\|Q_{V})$ and thus communication from the sensor cannot improve the Stein-exponent even for DMCs satisfying (15).
In contrast, we have the strict inequality
$\theta^{*}_{\textnormal{covert}}(\epsilon)>D(P_{V}\|Q_{V})$ (19)
for partially-connected DMCs whenever $T_{U}\neq P_{U}$ and for fully-connected DMCs whenever $T_{U}\in\bar{\mathcal{P}}_{U}(x_{1})$ for some $x_{1}$.
Proof:
We simply prove (19). To this end, start by noticing that for any $\pi_{UV}$ with marginal $\pi_{V}=P_{V}$:
$D\!\big(\pi_{UV}\big\|Q_{UV}\big)$ (20)
$=$ $D\!\big(P_{V}\big\|Q_{V}\big)+\sum_{v}P_{V}(v)\,D\!\big(\pi_{U|V}(\cdot|v)\big\|Q_{U|V}(\cdot|v)\big).$
The second summand is strictly positive, except when $\pi_{U|V}(u|v)=Q_{U|V}(u|v)$ for all $u$ and all $v$ with $P_{V}(v)>0$. Thus, to deduce that $E_{1}>D(P_{V}\|Q_{V})$ or $E_{2}>D(P_{V}\|Q_{V})$ it suffices to prove that $\pi_{U|V}=Q_{U|V}$ is not a valid choice in the minimizations. Inspecting the two minimizations, we see that $\pi_{U|V}=Q_{U|V}$ is not a permissible choice in the minimization of $E_{1}$ when $T_{U}\neq P_{U}$ and of $E_{2}$ when $T_{U}\in\ \bar{\mathcal{P}}_{U}(x_{1})$ for some $x_{1}$.
Notice also that $E_{3}>D(P_{V}\|Q_{V})$ by Assumptions (7). ∎
We now provide an example for which $\theta_{\textnormal{covert}}^{*}(\epsilon)>D(P_{V}\|Q_{V})$ even for fully-connected DMCs.
Example 1.
Let $P_{U}$ be Bern(0.2)(0.2) and $Q_{U}$ be Bern(0.7)(0.7). We assume degenerate observations V=V=const. at the receiver under both hypotheses.
Consider a BSC($0.4$) for both $\Gamma_{Z|X}$ and $\Gamma_{Y|X}$.
We start by characterizing the set $\bar{\mathcal{P}}_{U}(1)$. To this end, parametrize $\pi_{z}(0)=1-q$ and $\pi_{z}(1)=q$ and write
$)\displaystyle D(\pi_{z}\,\|\,\Gamma_{Z|X}(\cdot\mid 0))-\tfrac{3}{2}D(\pi_{z}\,\|\,\Gamma_{Z|X}(\cdot\mid\hat{x}$)) (21)
$=$ $(1-q)\log\frac{1-q}{0.6}+q\log\frac{q}{0.4}$
$-\ \frac{3}{2}\Big[(1-q)\log\frac{1-q}{0.4}+q\log\frac{q}{0.6}\Big]$
$=$ $\hskip-2.84544pt\frac{1}{2}H_{\textnormal{q}}(q)+\left(\frac{3}{2}-\frac{5}{2}q\right)\log 0.4+\left(-1+\frac{5}{2}q\right)\log 0.6.$ (22)
Above function is strictly concave in $q$ and thus the maximum is obtained by setting the derivative to 0:
$q^{*}=\frac{0.6^{5}}{0.4^{5}+0.6^{5}}=0.884.$ (23)
Plugging back this value into (22) and simplifying, we obtain:
$)\displaystyle\max_{\pi_{z}}D(\pi_{z}\,\|\,\Gamma_{Z|X}(\cdot\mid 0))-\tfrac{3}{2}D(\pi_{z}\,\|\,\Gamma_{Z|X}(\cdot\mid\hat{x}$)) (24)
$=$ $\frac{1}{2}\log\frac{0.4^{5}+0.6^{5}}{0.4^{2}0.6^{2}}=0.306\hskip 85.35826pt$
and thus
$\bar{\mathcal{P}}_{U}(1)=\left\{\pi_{U}\colon D(\pi_{U}\|P_{U})\geq\frac{1}{2}\log\frac{0.4^{5}+0.6^{5}}{0.4^{2}0.6^{2}}\right\}.$ (25)
Parameterizing the binary type
$\pi_{U}(1)=m,\quad\pi_{U}(0)=1-m,$ (26)
for $m\in[0,1]$, numerical evaluation shows that $\bar{\mathcal{P}}_{U}(1)$ contains all types $\pi_{U}$ parametrized by $m\geq 0.45$, i.e.,
$\bar{\mathcal{P}}_{U}(1)=\{\pi_{U}\colon\pi_{U}(1)\geq 0.45\}.$ (27)
We are now ready to evaluate the exponents $E_{1},E_{2},E_{3}$:
$E_{1}$ $=$ $D(P_{U}\|Q_{U})=0.7706$ (28)
$E_{2}$ $=$ $\min_{\pi_{U}\notin\bar{\mathcal{P}}_{U}(1)}D(\pi_{U}\|Q_{U})=0.2095$ (29)
$E_{3}$ $=$ $D(\Gamma_{Y|X}(\cdot|0)\|\Gamma_{Y|X}(\cdot|1))=0.1170.$ (30)
IV Proof of achievability results
To prove the achievability results in (16) and (17), we state the following lemma.
Lemma 1.
Assume that $X^{n}=0^{n}$ with probability $1-\delta_{n}$ and $X^{n}=x_{1}^{n}$ with probability $\delta_{n}>0$, then the warden’s divergence satisfies
$D\left(P_{Z^{n}}\,\big\|\,\Gamma_{Z|X}^{\otimes n}(\cdot|0^{n})\right)\leq\delta_{n}^{2}\chi^{2}\!\big(\Gamma_{Z|X}^{\otimes n}(\cdot|x_{1}^{n})\,\big\|\,\Gamma_{Z|X}^{\otimes n}(\cdot|0^{n})\big),$ (31)
for $\chi^{2}(\cdot\|\cdot)$ the chi-squared distance between two distributions:
$\chi^{2}(P\|Q):=\sum_{z\in\mathcal{Z}}\frac{\big(P(z)-Q(z)\big)^{2}}{Q(z)}.$
Moreover,
$\chi^{2}\!\big(\Gamma_{Z|X}^{\otimes n}(\cdot|x_{1}^{n})\,\big\|\,\Gamma_{Z|X}^{\otimes n}(\cdot|0^{n})\big)$ (32)
$\leq$ $(n+1)^{|\mathcal{Z}|}\;\max_{\begin{subarray}{c}\pi\in\mathcal{P}_{n}(\mathcal{Z})\end{subarray}}\;e^{-nD\!\big(\pi\big\|\Gamma_{Z|X}(\cdot|x_{1})\big)}\,$
$\cdot\ \Bigg(e^{-n\big(D(\pi\|\Gamma_{Z|X}(\cdot|x_{1}))-D(\pi\|\Gamma_{Z|X}(\cdot|0))\big)}-1\Bigg)^{2}.$
Proof:
See Appendix A. ∎
With this result at hand, we now prove the achievability results. Notice that in our proposed schemes, the sensor and decision center do not use their shared secret key $S$.
IV-A Achievability of (16)
Fix a small number $\mu>0$ and let $x_{1},y^{*}$ be as in the theorem.
Choose further a function $k(\cdot)$ satisfying
$\lim_{n\to\infty}k(n)$ $=$ $\infty$ (33a)
$\lim_{n\to\infty}\frac{k(n)}{n}$ $=$ $0.$ (33b)
We propose a scheme where the sensor uses only the first $k(n)$ channel uses, and sends 0 during the rest of the time.
For ease of notation we will also write $k$ instead of $k(n)$. The scheme works as follows
Sensor: If $U^{n}\in\mathcal{T}^{(n)}_{\mu}(P_{U})$, the sensor sends $X^{k}=0^{k}$. Otherwise, it sends $X^{k}=x_{1}^{k}$.
Decision Center: If at least one of the channel outputs is $y^{*}$ and if $V^{n}\in\mathcal{T}^{(n)}_{\mu}(P_{V})$, then it declares $\hat{\mathcal{H}}=0$. Otherwise, it declares $\hat{\mathcal{H}}=1$.
As shown in [16], the Type-I error probability $\alpha_{n}$ tends to 0 as $n\rightarrow\infty$ and $\mu\rightarrow 0$, while the Type-II error probability $\beta_{n}$ exceeds $E_{1}$. It remains to verify that the scheme satisfies the covertness constraint (11).
Analysis of the covertness constraint: Apply Lemma 1 but only to the first $k$ channel uses because the remaining channel uses do not contribute to the divergence. This allows to obtain:
$D\!\left(P_{Z^{n}|\mathcal{H}_{0}}\,\big\|\,\Gamma_{Z|X}^{\otimes n}(\cdot|0^{n})\right)$ (34)
$=$ $D\!\left(P_{Z^{k}|\mathcal{H}_{0}}\,\big\|\,\Gamma_{Z|X}^{\otimes k}(\cdot|0^{k})\right)$
$\leq$ $\delta_{n}^{2}\,\chi^{2}\!\big(\Gamma_{Z|X}^{\otimes k}(\cdot|x_{1}^{k})\,\big\|\,\Gamma_{Z|X}^{\otimes k}(\cdot|0^{k})\big)$ (36)
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $(k+1)^{|\mathcal{Z}|}2|\mathcal{U}|e^{-2n\mu^{2}}\max_{\begin{subarray}{c}\pi\in\mathcal{P}_{k}(\mathcal{Z})\end{subarray}}\ e^{-kD\!\big(\pi\big\|\Gamma_{Z|X}(\cdot|0)\big)}\,$
$\cdot\ \Bigg(e^{-k\big(D(\pi\|\Gamma_{Z|X}(\cdot|x_{1}))-D(\pi\|\Gamma_{Z|X}(\cdot|0))\big)}-1\Bigg)^{2},$
where in $(a)$ the second part of Lemma 1 and [21, Remark to Lemma 2.12].
Since $k$ grows only sublinearly in $n$ and $\mu>0$, we can conclude that the right-hand side of (36) tends to 0 almost exponentially fast in $n$, thus satisfying the covertness constraint (11).
Notice that our proposed scheme does not require the usage of the shared secret key $S$ between the transmitter and the receiver.
IV-B Achievability result in (17)
Pick an input symbol $x_{1}\in\mathcal{X}\setminus\{0\}$ and fix a small constant $\mu>0$.
We consider the following scheme:
Sensor: If $\pi_{U}\in\bar{\mathcal{P}}_{U}(x_{1})$, it transmits $X^{n}=x_{1}^{n}$. Otherwise, it transmits $X^{n}=0^{n}$.
Decision Center: If $V^{n}\in\mathcal{T}^{(n)}_{\mu}(P_{V})$ and $Y^{n}\in\mathcal{T}^{(n)}_{\mu}(\Gamma_{Y|X=0})$, it declares $\hat{\mathcal{H}}=0$. Otherwise, $\hat{\mathcal{H}}=1$.
We analyze above scheme.
Analysis of covertness constraint: We again employ Lemma 1, but now to the entire blocklength, and we again use the bound
$\delta_{n}:=\Pr[X^{n}=x_{1}^{n}]\leq(n+1)^{|\mathcal{U}|}\;\max_{\begin{subarray}{c}\pi_{U}\in\mathcal{\bar{P}}_{U}(x_{1})\end{subarray}}e^{-nD(\pi_{U}\|P_{U})}.$ (37)
Similarly to above, we obtain inequalities (41)-(41) on the top of the next page.
$D\left(P_{Z^{n}|\mathcal{H}=0}\,\big\|\,\Gamma_{Z|X}^{\otimes n}(\cdot|0^{n})\right)$ (41)
$\leq$ $\delta^{2}\,\chi^{2}\!\Big(\Gamma_{Z|X}^{\otimes n}(\cdot|x_{1}^{n})\,\big\|\,\Gamma_{Z|X}^{\otimes n}(\cdot|0^{n})\Big)$
$\leq$ $(n+1)^{2|\mathcal{U}|}\;\max_{\begin{subarray}{c}\pi_{U}\in\mathcal{\bar{P}}_{U}(x_{1})\end{subarray}}e^{-2nD(\pi_{U}\|P_{U})}\cdot(n+1)^{|\mathcal{Z}|}\max_{\begin{subarray}{c}\pi\in\mathcal{P}_{n}(\mathcal{Z})\end{subarray}}\Bigg\{\;e^{-nD(\pi\|\Gamma_{Z|X}(\cdot|x_{1}))}\;$
$\hskip 256.0748pt\Bigg(e^{-n\big(D(\pi\|\Gamma_{Z|X}(\cdot|x_{1}))-D(\pi\|\Gamma_{Z|X}(\cdot|0))\big)}-1\Bigg)^{2}\Bigg\}$
$=$ $(n+1)^{2|\mathcal{U}|}\;\max_{\begin{subarray}{c}\pi_{U}\in\mathcal{\bar{P}}_{U}(x_{1})\end{subarray}}(n+1)^{|\mathcal{Z}|}\;\max_{\begin{subarray}{c}\pi\in\mathcal{P}_{n}(\mathcal{Z})\end{subarray}}\;\Bigg\{e^{-n\big(3D(\pi\|\Gamma_{Z|X}(\cdot|x_{1}))-2D(\pi\|\Gamma_{Z|X}(\cdot|0))+2D(\pi_{U}\|Q_{U})\big)}$
$\hskip 73.97733pt-2e^{-n\big(2D(\pi\|\Gamma_{Z|X}(\cdot|x_{1}))-D(\pi\|\Gamma_{Z|X}(\cdot|0))+2D(\pi_{U}\|P_{U})\big)}+e^{-n\big(D\big(\pi\|\Gamma_{Z|X}(\cdot|x_{1})+2nD(\pi_{U}\|Q_{U})\big)\big)}\Bigg\}$
$\leq(n+1)^{2|\mathcal{U}|}\;(n+1)^{|\mathcal{Z}|}\max_{\begin{subarray}{c}\pi_{U}\in\mathcal{\bar{P}}_{U}(x_{1})\end{subarray}}\;\max_{\begin{subarray}{c}\pi\in\mathcal{P}_{n}(\mathcal{Z})\end{subarray}}\;\Bigg\{e^{-n\big(3D(\pi\|\Gamma_{Z|X}(\cdot|x_{1}))-2D(\pi\|\Gamma_{Z|X}(\cdot|0))+2D(\pi_{U}\|P_{U})\big)}$
$\hskip 85.35826pt-2e^{-n\big(2D(\pi\|\Gamma_{Z|X}(\cdot|x_{1}))-D(\pi\|\Gamma_{Z|X}(\cdot|0))+2D(\pi_{U}\|P_{U})\big)}+e^{-n\big(D(\pi\|\Gamma_{Z|X}(\cdot|x_{1}))+2D(\pi_{U}\|P_{U})\big)}\Bigg\}.$
As we argue next, all three exponential terms in (41) vanish as $n\to\infty$ because the terms multiplying n-n are positive. In fact,
$D(\pi\|\Gamma_{Z|X}(\cdot|x_{1}))+2D(\pi_{U}\|P_{U})>0$ (42)
because $P_{U}\notin\bar{\mathcal{P}}_{U}(x_{1})$ and thus the second divergence is strictly positive for all $\pi_{U}$ and $\pi$. Similarly, by the definition of the set $\bar{\mathcal{P}}_{U}(x_{1})$:
$3D(\pi\|\Gamma_{Z|X}(\cdot|x_{1}))-2D(\pi\|\Gamma_{Z|X}(\cdot|0))+2D(\pi_{U}\|P_{U})\big)>0.$
Finally, to see that also
$2D(\pi\|\Gamma_{Z|X}(\cdot|x_{1}))-D(\pi\|\Gamma_{Z|X}(\cdot|0))+2D(\pi_{U}\|P_{U})>0$
distinguish the cases where $D(\pi\|\Gamma_{Z|X}(\cdot|x_{1}))>D(\pi\|\Gamma_{Z|X}(\cdot|0))$ or not. In the former case, positivity of (IV-B) is obvious. In the latter case, positivity of (IV-B) implies also positivity of (IV-B), thus concluding that the divergence $d_{n}$ tends to 0 exponentially fast in the blocklength $n$.
We next examine the Type-I and Type-II error probabilities of our scheme.
Analysis of $\alpha_{n}$:
$1-\alpha_{n}$ (49)
$=$ $\Pr\left[\hat{\mathcal{H}}=0|\mathcal{H}=0\right]$
[…]
$\left.\hskip 71.13188ptY^{n}\in\mathcal{T}^{(n)}_{\mu}(\Gamma_{Y|X}(\cdot|0))|\mathcal{H}=0\right]$
$\geq$ $\Pr\left[V^{n}\in\mathcal{T}^{(n)}_{\mu}(P_{V})\textnormal{ and }Y^{n}\in\mathcal{T}^{(n)}_{\mu}(\Gamma_{Y|X}(\cdot|0))\right$.
 $\left.\hskip 91.04881pt\textnormal{ and }\pi_{U}\notin\mathcal{\bar{P}}_{U}(x_{1})|\mathcal{H}=0\right]$
$=$ $\Pr\left[V^{n}\in\mathcal{T}^{(n)}_{\mu}(P_{V})\textnormal{ and }\pi_{U}\notin\mathcal{\bar{P}}_{U}(x_{1})|\mathcal{H}=0\right]$
$\cdot\Pr\left[Y^{n}\in\mathcal{T}^{(n)}_{\mu}(\Gamma_{Y|X}(\cdot|0))|V^{n}\in\mathcal{T}^{(n)}_{\mu}(P_{V}),\right$.
$\left.\hskip 105.2751pt\pi_{U}\notin\mathcal{\bar{P}}_{U}(x_{1}),\mathcal{H}=0\right]$
$=$ $\Pr\left[V^{n}\in\mathcal{T}^{(n)}_{\mu}(P_{V})\textnormal{ and }\pi_{U}\notin\mathcal{\bar{P}}_{U}(x_{1})|\mathcal{H}=0\right]$
$\cdot\ \Pr\left[Y^{n}\in\mathcal{T}^{(n)}_{\mu}(\Gamma_{Y|X}(\cdot|0))|X^{n}=0^{n}\right]$
$\geq$ $\Pr\left[V^{n}\in\mathcal{T}^{(n)}_{\mu}(P_{V})\textnormal{ and }U^{n}\in\mathcal{T}^{(n)}_{\mu}(P_{U})|\mathcal{H}=0\right]$
$\ \ \cdot\ \Pr\left[Y^{n}\in\mathcal{T}^{(n)}_{\mu}(\Gamma_{Y|X}(\cdot|0))|X^{n}=0^{n}\right],$
where we used that $P_{U}$ and all surrounding types $\pi_{U}$ do not belong to $\bar{\mathcal{P}}_{U}(x_{1})$.
By the weak law of large numbers, both probabilities in (49) tend to 1 as $n\to\infty$ and thus
$\lim_{n\to\infty}\alpha_{n}=0.$ (50)
Analysis of $\beta_{n}$ and $\theta$:
$\beta_{n}$ $=$ $\Pr\left[\hat{\mathcal{H}}=0|\mathcal{H}=1\right]$ (57)
[…]
$\left.\hskip 56.9055ptY^{n}\in\mathcal{T}^{(n)}_{\mu}(\Gamma_{Y|X}(\cdot|0))|\mathcal{H}=1\right]$
$=$ $\Pr\left[V^{n}\in\mathcal{T}^{(n)}_{\mu}(P_{V}),\ \pi_{U}\notin\mathcal{\bar{P}}_{U},\ \right$.
$\left.\hskip 56.9055ptY^{n}\in\mathcal{T}^{(n)}_{\mu}(\Gamma_{Y|X}(\cdot|0))|\mathcal{H}=1\right]$
$+\ \Pr\left[V^{n}\in\mathcal{T}^{(n)}_{\mu}(P_{V}),\ \pi_{U}\in\mathcal{\bar{P}}_{U},\right$.
$\left.\hskip 56.9055pt\ Y^{n}\in\mathcal{T}^{(n)}_{\mu}(\Gamma_{Y|X}(\cdot|0))|\mathcal{H}=1\right]$
$=$ $\Pr\left[V^{n}\in\mathcal{T}^{(n)}_{\mu}(P_{V}),\pi_{U}\notin\mathcal{\bar{P}}_{U},\ \right$.
$\left.\hskip 54.06023ptY^{n}\in\mathcal{T}^{(n)}_{\mu}(\Gamma_{Y|X}(\cdot|0))|\mathcal{H}=1\right]$
$+\ \Pr\left[V^{n}\in\mathcal{T}^{(n)}_{\mu}(P_{V}),\ \pi_{U}\in\mathcal{\bar{P}}_{U}|\mathcal{H}=1\right]$
$\hskip 19.91684pt\cdot\ \Pr\left[Y^{n}\in\mathcal{T}^{(n)}_{\mu}(\Gamma_{Y|X}(\cdot|0))|X^{n}=x_{1}^{n}\right]$
$\leq$ $\Pr\left[V^{n}\in\mathcal{T}^{(n)}_{\mu}(P_{V}),\ \pi_{U}\notin\mathcal{\bar{P}}_{U}|\mathcal{H}=1\right]$
$+\ \Pr\left[V^{n}\in\mathcal{T}^{(n)}_{\mu}(P_{V}),\ \pi_{U}\in\mathcal{\bar{P}}_{U}|\mathcal{H}=1\right]$
$\hskip 11.38092pt\cdot\ \Pr\left[Y^{n}\in\mathcal{T}^{(n)}_{\mu}(\Gamma_{Y|X}(\cdot|0))|X^{n}=x_{1}^{n}\right]$
$=$ $\sum_{\begin{subarray}{c}\pi_{UV}:\\ \pi_{U}\notin\bar{\mathcal{P}}_{U}\\ \pi_{V}=P_{V}\end{subarray}}\Pr\!\left[\,\pi_{U^{n}V^{n}}=\pi_{UV}\,\middle|\,\mathcal{H}_{1}\right]\$
$+\sum_{\begin{subarray}{c}\pi_{UV}:\\ \pi_{U}\in\bar{\mathcal{P}}_{U}\\ \pi_{V}=P_{V}\end{subarray}}\Pr\!\left[\,\pi_{U^{n}V^{n}}=\pi_{UV}\,\middle|\,\mathcal{H}_{1}\right]$
$\hskip 19.91684pt\cdot\ \Pr\left[Y^{n}\in\mathcal{T}^{(n)}_{\mu}(\Gamma_{Y|X}(\cdot|0))|X^{n}=x_{1}^{n}\right]$
$\leq$ $(n+1)^{|\mathcal{U}||\mathcal{V}|}\max_{\begin{subarray}{c}\pi_{UV}:\\ \pi_{U}\notin\bar{\mathcal{P}}_{U}\\ \pi_{V}=P_{V}\end{subarray}}2^{-nD\left(\pi_{UV}\|Q_{UV}\right)}$
$+\ (n+1)^{|\mathcal{U}||\mathcal{V}|}(n+1)^{|\mathcal{Y}|}\ \max_{\begin{subarray}{c}\pi_{UV}:\\ \pi_{U}\in\bar{\mathcal{P}}_{U}\\ \pi_{V}=P_{V}\end{subarray}}$
$\hskip 11.38092pt2^{-n\left(D(\pi_{UV}\|Q_{UV})+\ D(\Gamma_{Y|X}(\cdot|0)\|\Gamma_{Y|X}(\cdot|x_{1}))\right)}.$
We can thus conclude that
[…]
This establishes achievability of the exponent in (17).
V Converse result in (16)
The converse is inspired by [16]. We again assume a stronger setup where the decision center directly observes the DMC inputs $X^{n}$ instead of the outputs $Y^{n}$. A converse result for this new setup implies also a converse for the original setup. Thus, we assume in the following that (8) be replaced by
$\hat{\mathcal{H}}=\tilde{g}^{(n)}(V^{n},X^{n})\in\{0,1\}.$ (59)
Before starting the converse proof for the new setup, we state some useful definitions and lemmas.
Define for each $n$:
$a_{n,i}^{0}$ $:=$ $\Pr[X_{i}\neq 0|\mathcal{H}=0],\quad i\in\{1,\ldots,n\},$ (60)
[…]
and
$\bar{a}_{n}:=\sqrt{\frac{1}{n}\sum_{i=1}^{n}a_{n,i}^{0}}.$ (62)
Define further the set of all low-weight inputs
$\tilde{\mathcal{X}}^{n}:=\{x^{n}\in\mathcal{X}^{n}\colon w_{\textnormal{H}}(x^{n})<\tilde{a}_{n}\cdot n\}.$ (63)
Lemma 2 (Covertness Constraint Implies Low-Weight Inputs).
For any sequence of encoding functions $\{f^{(n)}\}$ that satisfies the covertness constraint (11):
$\lim_{n\to\infty}\Pr\left[X^{n}\notin\tilde{\mathcal{X}}^{n}\,\big|\,\mathcal{H}=0\right]$ $=$ $1,$ (64)
$\lim_{n\to\infty}\frac{1}{n}\log\left|\tilde{\mathcal{X}}^{n}\right|$ $=$ $0.$ (65)
Proof:
See Appendix B. ∎
Lemma 3.
Consider any sequence of encoding and improved decision functions $\{f^{(n)},\tilde{g}^{(n)}\}$. Consider further a vanishing sequence $\phi_{n}\to 0$, as $n\to\infty$. If for each $n$ there exists an input sequence $\bar{x}^{n}\in\tilde{\mathcal{X}}^{n}$ and key $\bar{s}\in\mathcal{K}$ satisfying
$\Pr\left[\hat{\mathcal{H}}=0,X^{n}=\bar{x}^{n}\Big|\mathcal{H}=0,S=\bar{s}\right]\geq e^{-n\phi_{n}},$ (66)
then
$\hskip-22.76228pt\Pr[\hat{\mathcal{H}}=0,X^{n}=\bar{x}^{n}|\mathcal{H}=1,S=\bar{s}]$ (67)
$\geq$ $2^{-n(\min_{\pi_{UV}}D(\pi_{UV}\|Q_{UV})+\eta_{n})},\hskip 28.45274pt$
where the minimum is over all $\pi_{UV}$ with marginals $P_{U}$ and $P_{V}$ and $\eta_{n}$ is a vanishing sequence in $n$ that does not depend on $\bar{x}^{n}$ nor $\bar{s}$ but only on $P_{UV},Q_{UV}$ and the sequence $\phi_{n}$.
Proof:
See Appendix C. ∎
We are now ready to prove the converse result. Fix any sequence of encoding and modified decision functions $\{f^{(n)},\tilde{g}^{(n)}\}$ satisfying the Type-I error constraint (12a) and the covertness constraint (11). Let $\tilde{\mathcal{X}}^{n}$ be defined as in (63).
Notice then that:
$\Pr\left[\hat{\mathcal{H}}=0,X^{n}\in\tilde{\mathcal{X}}^{n}|\mathcal{H}=0\right]$ (69)
$=$ $\Pr\left[\hat{\mathcal{H}}=0|\mathcal{H}=0\right]-\Pr\left[\hat{\mathcal{H}}=0,X^{n}\notin\tilde{\mathcal{X}}^{n}|\mathcal{H}=0\right]$
$\geq$ $1-\epsilon-\Pr[X^{n}\notin\tilde{\mathcal{X}}^{n}|\mathcal{H}=0].$
By Lemma 2 and because the chosen sequence of encoding functions satisfies the covertness constraint (11), for any $\eta\in(0,1-\epsilon)$ and sufficiently large $n$:
$\Pr\left[\hat{\mathcal{H}}=0,X^{n}\in\tilde{\mathcal{X}}^{n}|\mathcal{H}=0\right]\geq 1-\epsilon-\eta.$ (70)
In particular, there must be a special sequence $\bar{x}^{n}\in\tilde{\mathcal{X}}^{n}$ so that:
$\Pr\left[\hat{\mathcal{H}}=0,X^{n}=\bar{x}^{n}\Big|\mathcal{H}=0\right]\geq\frac{1-\epsilon-\eta}{|\tilde{\mathcal{X}}^{n}|}.$ (71)
Fix this sequence $\bar{x}^{n}$ and choose an arbitrary constant $c>1$. Define then the subset $\mathcal{\bar{S}}_{\bar{x}^{n}}\subseteq\mathcal{K}$
$\mathcal{\bar{S}}_{\bar{x}^{n}}$ (72)
[…]
$\Pr\left[\hat{\mathcal{H}}=0,X^{n}=\bar{x}^{n}\Big|\mathcal{H}=0,S=\bar{s}\right]\geq\frac{1-\epsilon-\eta}{c|\tilde{\mathcal{X}}^{n}|}\bigg\}.$
Notice that by Lemma 2 the cardinality of $|\tilde{\mathcal{X}}^{n}|$ grows subexponentially in $n$ and thus the ratio $\frac{1-\epsilon-\eta}{c|\tilde{\mathcal{X}}^{n}|}$ is of the form $e^{-n\phi_{n}}$ for a vanishing sequence $\phi_{n}$. I.e., for any $\bar{s}\in\mathcal{\bar{S}}_{\bar{x}^{n}}$:
$\Pr\left[\hat{\mathcal{H}}=0,X^{n}=\bar{x}^{n}\Big|\mathcal{H}=0,S=\bar{s}\right]\geq e^{-n\phi_{n}},$ (73)
for
$\phi_{n}:=\log\frac{c|\tilde{\mathcal{X}}^{n}|}{1-\epsilon-\eta}.$ (74)
Therefore, Lemma 3 implies that there exists a vanishing sequence $\eta_{n}$ so that for any $\bar{s}\in\mathcal{\bar{S}}_{\bar{x}^{n}}$:
$\Pr[\hat{\mathcal{H}}=0|\mathcal{H}=1,S=\bar{s}]\geq 2^{-n(\min_{\pi_{UV}}D(\pi_{UV}\|Q_{UV})+\eta_{n})}.$ (75)
As a consequence:
$\beta_{n}$ $=$ $\Pr\left[\hat{\mathcal{H}}=0|\mathcal{H}=1\right]$ (76)
$\geq$ $\Pr\left[\hat{\mathcal{H}}=0,X^{n}=\bar{x}^{n}|\mathcal{H}=1\right]$ (78)
$\geq$ $\sum_{s\in\mathcal{\bar{S}}_{\bar{x}^{n}}}\Pr[S=s]$
$\hskip 19.91684pt\cdot\ \Pr\left[\hat{\mathcal{H}}=0,X^{n}=\bar{x}^{n}|\mathcal{H}=1,S=s\right]$
$\geq$ $\Pr[S\in\mathcal{\bar{S}}_{\bar{x}^{n}}]\left(2^{-n(\min_{\pi_{UV}}D(\pi_{UV}\|Q_{UV})+\eta_{n})}\right),$ (79)
where the minimum is over all types $\pi_{UV}$ with marginals $P_{U}$ and $P_{V}$.
We next bound the probability of the set $\bar{s}\in\mathcal{\bar{S}}_{\bar{x}^{n}}$. To this end, notice
$\Pr\left[\hat{\mathcal{H}}=0,X^{n}=\bar{x}^{n}\Big|\mathcal{H}=0\right]$ (81)
$=$ $\sum_{s\in\mathcal{\bar{S}}_{\bar{x}^{n}}}\Pr\left[\hat{\mathcal{H}}=0,X^{n}=x^{n}|\mathcal{H}=0,S=s\right]\Pr[S=s]$
$+\ \sum_{s\notin\mathcal{\bar{S}}_{\bar{x}^{n}}}\Pr\left[\hat{\mathcal{H}}=0,X^{n}=x^{n}|\mathcal{H}=0,S=s\right]$
$\cdot\ \Pr[S=s]$
$\leq$ $\Pr\left[S\in\mathcal{\bar{S}}_{\bar{x}^{n}}\right]+\sum_{s\notin\bar{\mathcal{S}}_{\bar{x}^{n}}}\Pr\left[\hat{\mathcal{H}}=0,X^{n}=x^{n}|\mathcal{H}=0,S=s\right]$
$\cdot\ \Pr(S=s)$
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $\Pr\left[S\in\bar{\mathcal{S}}_{\bar{x}^{n}}\right]+\frac{1-\epsilon-\eta}{c|\tilde{\mathcal{X}}^{n}|}\bigg(1-\Pr\left[s\in\bar{\mathcal{S}}_{\bar{x}^{n}}\right]\bigg)$ (82)
$=$ $\Pr\left[S\in\bar{\mathcal{S}}_{\bar{x}^{n}}\right]Bigg(1-\frac{1-\epsilon-\eta}{c|\tilde{\mathcal{X}}^{n}|}\Bigg)+\frac{1-\epsilon-\eta}{c|\tilde{\mathcal{X}}^{n}|},$ (83)
where $(a)$ holds by the definition of $\bar{S}_{\bar{x}^{n}}$. Combining (71) and (83), we then obtain:
$\Pr\left[S\in\bar{\mathcal{S}}_{\bar{x}^{n}}\right]$ (84)
$\geq$ $\frac{1-\epsilon-\eta}{|\tilde{\mathcal{X}}^{n}|}\bigg(1-\frac{1}{c}\bigg)\Bigg(1-\frac{1-\epsilon-\eta}{c|\tilde{\mathcal{X}}^{n}|}\Bigg)^{-1}$
$\geq$ $\frac{1-\epsilon-\eta}{|\tilde{\mathcal{X}}^{n}|}\bigg(1-\frac{1}{c}\bigg).$ (85)
The proof is concluded by combining (79) with (85) and (65) amd by letting $n\to\infty$:
[…]
$=$ $\min_{\pi_{UV}}D(\pi_{UV}\|Q_{UV}).$ (88)
VI Conclusion
We studied distributed hypothesis testing under a covertness constraint in the non-alert setting, where an external warden must be unable to detect whether communication is taking place under the null hypothesis. For some partially-connected DMCs, we showed that the Stein exponent coincides with that of the noiseless sublinear-rate communication model of Shalaby and Papamarcou, and is independent of the specific transition law. For the remaining DMCs, we proposed an achievable Stein-exponent and showed that covert communication can strictly improve over the local exponent at the decision center. All proposed schemes operate without any shared secret key and satisfy a strong divergence-based covertness constraint. An interesting direction for future work is to establish a converse for fully connected DMCs case and to study covertness constraints imposed under both hypotheses.
Acknowledgments
The work of the authors has been supported by the ERC under grant agreement 101125691.
Appendix A Proof of Lemma 1
For the random tuple $Z^{n}\sim P_{Z^{n}|\mathcal{H}=0}$, we have:
$D\!\left(P_{Z^{n}|\mathcal{H}=0}\,\big\|\,\Gamma_{Z|X}^{\otimes n}(\cdot|0^{n})\right)$ (93)
${=}$ $D\!\Big((1-\delta_{n})\,\Gamma_{Z|X}^{\otimes n}(\,\cdot\,|0^{n})$
$\hskip 56.9055pt+\ \delta_{n}\,\Gamma_{Z|X}^{\otimes n}(\,\cdot\,|x_{1}^{n})\,\Big\|\,\Gamma_{Z|X}^{\otimes n}(\,\cdot\,|0^{n})\Big)$
$=$ $\sum_{z^{n}\in\mathcal{Z}^{n}}\Big[(1-\delta_{n})\,\Gamma_{Z|X}^{\otimes n}(z^{n}|0^{n})+\delta_{n}\,\Gamma_{Z|X}^{\otimes n}(z^{n}|x_{1}^{n})\Big]$
$\hskip 8.53581pt\cdot\ \log\frac{(1-\delta_{n})\,\Gamma_{Z|X}^{\otimes n}(z^{n}|0^{n})+p\,\Gamma_{Z|X}^{\otimes n}(z^{n}|x_{1}^{n})}{\Gamma_{Z|X}^{\otimes n}(z^{n}|0^{n})}$
$=$ $\sum_{z^{n}\in\mathcal{Z}^{n}}\Big[(1-\delta_{n})\,\Gamma_{Z|X}^{\otimes n}(z^{n}|0^{n})+\delta_{n}\,\Gamma_{Z|X}^{\otimes n}(z^{n}|x_{1}^{n})\Big]$
$\cdot\ \log\Bigg[1+\delta_{n}\,\frac{\Gamma_{Z|X}^{\otimes n}(z^{n}|x_{1}^{n})-\Gamma_{Z|X}^{\otimes n}(z^{n}|0^{n})}{\Gamma_{Z|X}^{\otimes n}(z^{n}|0^{n})}\Bigg]$
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $\sum_{z^{n}\in\mathcal{Z}^{n}}\Big[(1-\delta_{n})\,\Gamma_{Z|X}^{\otimes n}(z^{n}|0^{n})+\delta_{n}\,\Gamma_{Z|X}^{\otimes n}(z^{n}|x_{1}^{n})\Big]$
$\hskip 28.45274pt\cdot\ \delta_{n}\,\frac{\Gamma_{Z|X}^{\otimes n}(z^{n}|x_{1}^{n})-\Gamma_{Z|X}^{\otimes n}(z^{n}|0^{n})}{\Gamma_{Z|X}^{\otimes n}(z^{n}|0^{n})}$
$=$ $\delta_{n}\sum_{z^{n}\in\mathcal{Z}^{n}}(\Gamma_{Z|X}^{\otimes n}(z^{n}|x_{1}^{n})-\Gamma_{Z|X}^{\otimes n}(z^{n}|0^{n}))$
$\hskip 11.38092pt+\ \delta_{n}^{2}\sum_{z^{n}\mathcal{Z}^{n}}\frac{(\Gamma_{Z|X}^{\otimes n}(z^{n}|x_{1}^{n})-\Gamma_{Z|X}^{\otimes n}(z^{n}|0^{n}))^{2}}{\Gamma_{Z|X}^{\otimes n}(z^{n}|0^{n})}$
$=$ $\delta_{n}^{2}\sum_{z^{n}\in\mathcal{Z}^{n}}\frac{\big(\Gamma_{Z|X}^{\otimes n}(z^{n}|x_{1}^{n})-\Gamma_{Z|X}^{\otimes n}(z^{n}|0^{n})\big)^{2}}{\Gamma_{Z|X}^{\otimes n}(z^{n}|0^{n})}$ (94)
$=$ $\delta_{n}^{2}\,\chi^{2}\big(\Gamma_{Z|X}^{\otimes n}(\cdot|x_{1}^{n})\,\|\,\Gamma_{Z|X}^{\otimes n}(\cdot|0^{n})\big),$ (95)
where $(a)$ holds because $\log(1+x)\leq x$.
We now develop the above $\chi^{2}$-distance:
$\chi^{2}\!\big(\Gamma_{Z|X}^{\otimes n}(\cdot|x_{1}^{n})\,\big\|\,\Gamma_{Z|X}^{\otimes n}(\cdot|0^{n})\big)$ (98)
$=$ $\sum_{z^{n}\in\mathcal{Z}^{n}}\Gamma_{Z|X}^{\otimes n}(z^{n}|0^{n})\Bigg(\frac{\Gamma_{Z|X}^{\otimes n}(z^{n}|x_{1}^{n})}{\Gamma_{Z|X}^{\otimes n}(z^{n}|0)}-1\Bigg)^{2}$
$=$ $\sum_{\pi\in\mathcal{P}_{n}(\mathcal{Z})}\mathbb{P}_{Z^{n}\sim\Gamma_{Z|X}^{\otimes n}(\cdot|0)}\big(Z^{n}\in\mathcal{T}_{\pi}\big)$
$\hskip 28.45274pt\cdot\Bigg(\prod_{a\in\mathcal{Z}}\Bigg(\frac{\Gamma_{Z|X}(a|x_{1})}{\Gamma_{Z|X}(a|0)}\Bigg)^{\,n\pi(a)}-1\Bigg)^{2}$
$\leq$ $(n+1)^{|\mathcal{Z}|}\;\max_{\begin{subarray}{c}\pi\in\mathcal{P}_{n}(\mathcal{Z})\end{subarray}}\;e^{-nD\!\big(\pi\big\|\Gamma_{Z|X}(\cdot|0)\big)}\,$
$\cdot\Bigg(e^{-k\big(D(\pi\|\Gamma_{Z|X}(\cdot|x_{1}))-D(\pi\|\Gamma_{Z|X}(\cdot|0))\big)}-1\Bigg)^{2},$
where in (98), we used that
$\prod_{a\in\mathcal{Z}}\Bigg(\frac{\Gamma_{Z|X}(a|x_{1})}{\Gamma_{Z|X}(a|0)}\Bigg)^{\,n\pi(a)}$ (99)
$=$ $e^{-n\big(D(\pi\|\Gamma_{Z|X}(\cdot|x_{1}))-D(\pi\|\Gamma_{Z|X}(\cdot|0))\big)},$
which ends the proof.
Appendix B Proof of Lemma 2
Start by noticing that:
$D(P_{Z^{n}|\mathcal{H}=0}\|\Gamma_{Z|X}^{\otimes n}(\cdot|0^{n}))$ (100)
$=$ $-H(Z^{n})+\mathbb{E}_{Z^{n}}\left[\log\left(\frac{1}{\Gamma_{Z|X}^{\otimes n}(Z^{n}|0^{n})}\right)\right]$
$\overset{(a)}{\geq}$ $-\sum_{i=1}^{n}H(Z_{i})+\mathbb{E}_{Z_{i}}\left[\log\left(\frac{1}{\Gamma_{Z|X}(Z_{i}|0)}\right)\right]$ (101)
$=$ $\sum_{i=1}^{n}D\!\left(P_{Z_{i}|\mathcal{H}=0}\middle\|\Gamma_{Z|X}(\cdot\mid 0)\right)$ (103)
$\overset{(b)}{=}$ $\sum_{i=1}^{n}D\Big(\sum_{x\in\mathcal{X}\setminus\{0\}}b_{x}a_{n,i}^{0}\Gamma_{Z|X}(\cdot\mid x)$
$\hskip 14.22636pt+\ (1-a_{n,i}^{0})\Gamma_{Z|X}(\cdot\mid 0)\|\hskip 2.84544pt\Gamma_{Z|X}(\cdot\mid 0)\Big)$
where $(a)$ holds because conditioning reduces entropy; and $(b)$ because :
$P_{Z_{i}|\mathcal{H}=0}(z)$ (106)
$=$ $\Pr[Z_{i}=z\mid X_{i}=0,\mathcal{H}=0]\Pr[X_{i}=0\mid\mathcal{H}=0]$
$+\ \Pr[Z_{i}=z\mid X_{i}\neq 0,\mathcal{H}=0]\Pr[X_{i}\neq 0\mid\mathcal{H}=0]$
$=$ $(1-\Pr[X_{i}\neq 0\mid\mathcal{H}=0])$
$\cdot\ \Pr(Z_{i}=z\mid X_{i}=0,\mathcal{H}=0)+\Pr[X_{i}\neq 0\mid\mathcal{H}=0]$
$\cdot\ \sum_{x\in\mathcal{X}\setminus\{0\}}\Pr(Z_{i}=z\mid X_{i}=x,\mathcal{H}=0)$
$\cdot\ \Pr(X_{i}=x\mid X_{i}\neq 0,\mathcal{H}=0)$
$=$ $(1-a_{n,i}^{0})\,\Gamma_{Z|X}(z\mid 0)\ \$
$+\ a_{n,i}^{0}\sum_{x\in\mathcal{X}\setminus\{0\}}b_{x}\,\Gamma_{Z|X}(z\mid x).$
By the covertness constraint (11) for $\mathcal{H}=0$ and the nonnegativity of Kullback-Leibler divergence, we deduce that each summand on the right-hand side of (103) must vanish, and thus
$\lim_{n\to\infty}a_{n,i}^{0}=0,\quad i\in\{1,\ldots,n\},$ (107)
and
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}a_{n,i}^{0}=0.$ (108)
Notice that by Markov’s inequality:
$\sum_{i=1}^{n}a_{n,i}^{0}$ $=$ $\mathbb{E}[w_{\textnormal{H}}(X^{n})|\mathcal{H}=0]$ (109)
$\geq$ $\bar{a}_{n}\cdot n\cdot\Pr[X^{n}\notin\tilde{\mathcal{X}}^{n}|\mathcal{H}=0].$ (110)
By (62) this further implies
$\Pr\left[X^{n}\notin\tilde{\mathcal{X}}^{n}|\mathcal{H}=0\right]\leq\frac{\frac{1}{n}\sum_{i=1}^{n}a_{n,i}^{0}}{\bar{a}_{n}}$ $=$ $\sqrt{\frac{1}{n}\sum_{i=1}^{n}a_{n,i}^{0}}.$ (111)
By (108) this proves (64).
To see that the size of $\tilde{\mathcal{X}}^{n}$ does not grow exponentially, we notice that this set can be described as the union over all type-classes (i.e., sets of sequences with same type) for types that assign frequency larger or equal to $1-\bar{a}_{n}$ to the 0 symbol. Since the type-class for type $\boldsymbol{\pi}$ is of size at most $2^{nH(\boldsymbol{\pi})}$ and since the number of type-classes is bounded by $(n+1)^{|\mathcal{X}|}$, we have:
$|\tilde{\mathcal{X}}^{n}|$ $\leq$ $(n+1)^{|\mathcal{X}|}2^{n\max_{\boldsymbol{\pi}}H(\boldsymbol{\pi})},$ (112)
where the maximum is over all types $\boldsymbol{\pi}$ with $\boldsymbol{\pi}(0)\geq 1-\bar{a}_{n}$. Since $\bar{a}_{n}$ vanishes as $n\to\infty$ and by the continuity of the entropy functional, we obtain
$\lim_{n\to\infty}\frac{1}{n}\log\left|\tilde{\mathcal{X}}^{n}\right|$ $\leq$ $\lim_{n\to\infty}\left[\frac{|\mathcal{X}|}{n}\log(n+1)+\max_{\begin{subarray}{c}\boldsymbol{\pi}\colon\\ \boldsymbol{\pi}(0)\geq 1-\bar{a}_{n}\end{subarray}}H(\boldsymbol{\pi})\right]$ (113)
$=$ $0.$
Appendix C Proof of Lemma 3
Define the sets
$\bar{\mathcal{C}}_{\bar{s}}$ $:=$ $\{u^{n}\in\mathcal{U}^{n}\colon f^{(n)}(u^{n},\bar{s})=\bar{x}^{n}\}$ (114)
$\bar{\mathcal{D}}_{\bar{s}}$ $:=$ $\{v^{n}\in\mathcal{V}^{n}\colon g^{(n)}(v^{n},\bar{x}^{n},\bar{s})=0\}.$ (115)
Note that
$P_{UV}^{\otimes n}(\bar{\mathcal{C}}_{\bar{s}}\times\bar{\mathcal{D}}_{\bar{s}})$ (116)
$=$ $\hskip-17.07182pt\sum_{\begin{subarray}{c}v^{n},u^{n}\colon\\ f^{(n)}(u^{n},\bar{s})=\bar{x}^{n}\\ g^{(n)}(v^{n},\bar{x}^{n},\bar{s})=0\end{subarray}}\hskip-14.22636pt\Pr[U^{n}=u^{n},V^{n}=v^{n}|\mathcal{H}=0,S=\bar{s}]$
[…]
which by assumption exceeds $e^{-n\phi_{n}}$.
Pick now an arbitrary $\tilde{P}_{UV}$ with marginals $\tilde{P}_{U}=P_{U}$ and $\tilde{P}_{V}=P_{V}$ and notice
$\tilde{P}_{U}^{\otimes n}(\bar{\mathcal{C}}_{\bar{s}})$ $=$ $P_{U}^{\otimes n}(\bar{\mathcal{C}}_{\bar{s}})\geq e^{-n\phi_{n}}$ (118a)
and
$\tilde{P}_{V}^{\otimes n}(\bar{\mathcal{D}}_{\bar{s}})$ $=$ $P_{V}^{\otimes n}(\bar{\mathcal{D}}_{\bar{s}})\geq e^{-n\phi_{n}}.$ (118b)
In the following we shall slightly blow up (enlarge) the sets $\bar{\mathcal{C}}_{\bar{s}}$ and $\bar{\mathcal{D}}_{\bar{s}}$ to obtain $\hat{\mathcal{C}}_{\bar{s}}$ and $\hat{\mathcal{D}}_{\bar{s}}$, and then show that these enlarged sets contain a large portion of the typical set $\mathcal{T}_{\mu_{n}}^{(n)}(\tilde{P}_{UV})$. To this end, let $\{\ell_{n}\}_{n\geq 1}$ be a sequence satisfying $\lim_{n\to\infty}\ell_{n}/\sqrt{n}=\infty$ and $\lim_{n\to\infty}\ell_{n}/n=0$, and define the blown up regions
$\hat{\mathcal{C}}_{\bar{s}}$ $:=$ $\left\{\tilde{u}^{n}\colon\exists u^{n}\in\bar{\mathcal{C}}_{\bar{s}}\;\;{s.t.}\;\;d_{\text{H}}(\tilde{u}^{n},u^{n})\leq\ell_{n}\right\}$ (119)
$\hat{\mathcal{D}}_{\bar{s}}$ $:=$ $\left\{\tilde{v}^{n}\colon\exists v^{n}\in\bar{\mathcal{D}}_{\bar{s}}\;\;{s.t.}\;\;d_{\text{H}}(\tilde{v}^{n},v^{n})\leq\ell_{n}\right\}.$ (120)
By (118) and the blowing-up lemma [22, Remark on p. 446]:
$\tilde{P}_{U}^{\otimes n}(\hat{\mathcal{C}}_{\bar{s}})$ $\geq$ $1-\lambda_{n},$ (121a)
$\tilde{P}_{V}^{\otimes n}(\hat{\mathcal{D}}_{\bar{s}})$ $\geq$ $1-\lambda_{n},$ (121b)
for some sequence $\lambda_{n}$ that tends to 0 as $n\to\infty$.
Continue with the following considerations:
$\hskip-22.76228pt\Pr\left[\hat{\mathcal{H}}=0,X^{n}=\bar{x}^{n}|\mathcal{H}=1,S=\bar{s}\right]$ (122)
$=$ $Q_{UV}^{\otimes n}\left(\bar{\mathcal{C}}_{s}\times\bar{\mathcal{D}}_{s}\right)$
$\geq$ $Q_{UV}^{\otimes n}\left(\hat{\mathcal{C}}_{s}\times\hat{\mathcal{D}}_{s}\right)\cdot 2^{-n\xi_{n}},$ (123)
where
$\xi_{n}:=H_{\textnormal{b}}(\ell_{n}/n)+\frac{\ell_{n}}{n}\log(|\mathcal{U}||\mathcal{V}|)-\frac{\ell_{n}}{n}\log\min_{(u,v)}\underbrace{Q_{UV}(u,v)}_{>0},$ (124)
and the last inequality is obtained by simple counting arguments and because $2^{nH_{\textnormal{b}}(\ell_{n}/n)}$ upper bounds the set of all binary vectors with Hamming weight $\ell_{n}/n$.
Let $\{\mu_{n}\}_{n=1}^{\infty}$ denote a sequence of positive numbers satisfying [^1] Condition (125b) ensures that the probability of the strongly typical set $\mathcal{T}_{\mu_{n}}^{(n)}(P_{XY})$ under $P_{XY}^{\otimes n}$ converges to one as $n\to\infty$; see [21, Remark following Lemma 2.12].
$\lim_{n\to\infty}\mu_{n}$ $=$ $0,$ (125a)
$\lim_{n\to\infty}n\mu_{n}^{2}$ $=$ $\infty.$ (125b)
Define the new set:
$\hat{\mathcal{E}}_{\bar{s}}\triangleq\{(u^{n},v^{n})\in\mathcal{T}_{\mu_{n}}^{(n)}(\tilde{P}_{UV})\colon u^{n}\in\hat{\mathcal{C}}_{\bar{s}},\;v^{n}\in\hat{\mathcal{D}}_{\bar{s}}\}.$ (126)
Denote the probability of this set under $P_{UV}^{\otimes}$ by $\Delta_{\bar{s}}$,
$\Delta_{\bar{s}}:=P_{UV}^{\otimes n}\left(\hat{\mathcal{E}}_{\bar{s}}\right),$ (127)
and define the tuple $(\tilde{U}^{n},\tilde{V}^{n})$ to be of joint pmf
$P_{\tilde{U}^{n}\tilde{V}^{n}}(u^{n},{v}^{n})=\frac{P_{{UV}}^{\otimes n}(u^{n},{v}^{n})}{\Delta_{\bar{s}}}\cdot\mathbbm{1}\left\{(u^{n},v^{n})\in\hat{\mathcal{E}}_{\bar{s}}\right\}.$ (128)
We continue to bound the exponent in (123). Since $\hat{\mathcal{E}}_{\bar{s}}$ is a subset of $\hat{\mathcal{C}}_{\bar{s}}\times\hat{\mathcal{D}}_{\bar{s}}$, we have:
$-\frac{1}{n}\log\Pr\left[\hat{\mathcal{H}}=0|\mathcal{H}=1,S=\bar{s}\right]$ (129)
$\leq$ $-\frac{1}{n}\log\Pr\left[\hat{\mathcal{H}}=0,X^{n}=\bar{x}^{n}|\mathcal{H}=1,S=\bar{s}\right]$
$=$ $-\frac{1}{n}\log\bigg(Q_{UV}^{\otimes n}\left(\hat{\mathcal{C}}_{\bar{s}}\times\hat{\mathcal{D}}_{\bar{s}}\right)\cdot 2^{-n\xi_{n}}\bigg)$ (130)
$\leq$ $-\frac{1}{n}\log\bigg(Q_{UV}^{\otimes n}\left(\hat{\mathcal{E}}_{\bar{s}}\right)\bigg)+\xi_{n}$ (131)
$\stackrel{{\scriptstyle(a)}}{{=}}$ $\frac{1}{n}P_{\tilde{U}^{n}\tilde{V}^{n}}\left(\hat{\mathcal{E}}_{\bar{s}}\right)\log\frac{P_{\tilde{U}^{n}\tilde{V}^{n}}\left(\hat{\mathcal{E}}_{\bar{s}}\right)}{Q_{UV}^{\otimes n}\left(\hat{\mathcal{E}}_{\bar{s}}\right)}+\xi_{n}$ (132)
$\stackrel{{\scriptstyle(b)}}{{\leq}}$ $\frac{1}{n}D\left(P_{\tilde{U}^{n}\tilde{V}^{n}}\|Q_{UV}^{\otimes n}\right)+\xi_{n}$ (137)
$\stackrel{{\scriptstyle(c)}}{{=}}$ $\frac{1}{n}\sum_{(u^{n},v^{n})\in\hat{\mathcal{E}}_{\bar{s}}}P_{\tilde{U}^{n}\tilde{V}^{n}}(u^{n},v^{n})\log\frac{P_{UV}^{\otimes n}(u^{n},v^{n})}{Q_{UV}^{\otimes n}(u^{n},v^{n})}$
$\hskip 8.53581pt-\frac{1}{n}\log\Delta_{\bar{s}}+\xi_{n}$
$\leq$ $\frac{1}{n}\sum_{i=1}^{n}\sum_{u_{i},v_{i}}P_{\tilde{U}_{i}\tilde{V}_{i}}(u_{i},v_{i})\log\frac{P_{UV}(u_{i},v_{i})}{Q_{UV}(u_{i},v_{i})}$
$\hskip 14.22636pt-\frac{1}{n}\log\Delta_{\bar{s}}+\xi_{n}$
$=$ $\sum_{u,v}P_{\tilde{U}_{T}\tilde{V}_{T}}(u,v)\log\frac{P_{UV}(u,v)}{Q_{UV}(u,v)}$
$-\frac{1}{n}\log\Delta_{\bar{s}}+\xi_{n}$
$=$ $\sum_{u,v}\tilde{P}_{UV}(u,v)\log\frac{P_{UV}(u,v)}{Q_{UV}(u,v)}$
$-\frac{1}{n}\log\Delta_{\bar{s}}+o(1),$
where $T$ is a uniform random variable over $\{1,\ldots,n\}$ independent of $(\tilde{U}^{n},\tilde{V}^{n})$ and moreover $o(1)$ is a function that tends to 0 as $n\to\infty$ and does not depend on neither $\bar{x}^{n}$ nor $\bar{s}$.
In above, $(a)$ holds because $P_{\tilde{U}^{n}\tilde{V}^{n}}$ is only defined on $\hat{\mathcal{E}}_{\bar{s}}$ and thus $P_{\tilde{U}^{n}\tilde{V}^{n}}(\hat{\mathcal{E}}_{\bar{s}})=1$; $(b)$ follows from the data processing inequality, since (132) is the KL divergence between the two binary distributions obtained by applying $P_{\tilde{U}^{n}\tilde{V}^{n}}$ and $Q_{UV}^{\otimes n}$ to the indicator function $\mathbbm{1}\{(u^{n},v^{n})\in\mathcal{E}_{\bar{s}}\}$ and $(c)$ follows from the definition of $P_{\tilde{U}^{n}\tilde{V}^{n}}$.
We shall conclude the desired proof by showing below that
$\Delta_{\bar{s}}$ (138)
$\geq$ $\left(1-2\lambda_{n}-\frac{|\mathcal{U}||\mathcal{V}|}{4\mu_{n}^{2}n}\right)$
$\cdot\ 2^{-\max_{\hat{\boldsymbol{\pi}}_{UV}}nD(\hat{\boldsymbol{\pi}}_{UV}\|P_{UV})}\frac{1}{(n+1)^{|\mathcal{U}||\mathcal{V}|}}$
where the maximization is over all types $\hat{\boldsymbol{\pi}}_{UV}$ that are of distance less than $\mu_{n}$ from $\tilde{P}_{UV}$. In other words,
$\Delta_{\bar{s}}\geq 2^{-n(D(\tilde{P}_{UV}\|P_{UV})+o(1))}$ (139)
where the $o(1)$ function again does not depend on neither $\bar{x}^{n}$ nor $\bar{s}$.
Combined with (137), this establishes the desired statement in the lemma. In fact, notice that
$\hskip-85.35826pt\sum_{u,v}\tilde{P}_{UV}(u,v)\log\frac{P_{UV}(u,v)}{Q_{UV}(u,v)}+D(\tilde{P}_{UV}\|P_{UV})$ (140)
$\hskip-56.9055pt=D(\tilde{P}_{UV}\|Q_{UV}).$
It remains to show (138), where we proceed similarly to [10]. By the union bound and by (121) and standard bounds on the probability of the typical set[21]
$1-\tilde{P}_{UV}^{\otimes n}(\hat{\mathcal{E}}_{\bar{s}})$ (142)
$=$ $\Pr\left[{U}^{n}\notin\hat{\mathcal{C}}_{\bar{s}}\textnormal{ or }{V}^{n}\notin\hat{\mathcal{D}}_{\bar{s}}\right$.
 $\hskip 19.91692pt\left.\textnormal{ or }(U^{n},V^{n})\notin\mathcal{T}_{\mu_{n}}^{(n)}(\tilde{P}_{UV})|\mathcal{H}=0\right]$
$\leq$ $\Pr\left[{U}^{n}\notin\hat{\mathcal{C}}_{\bar{s}}|\mathcal{H}=0\right]+\Pr\left[{V}^{n}\notin\hat{\mathcal{D}}_{\bar{s}}|\mathcal{H}=0\right]\ \$
$+\ \Pr\left[(U^{n},V^{n})\notin\mathcal{T}_{\mu_{n}}^{(n)}(\tilde{P}_{UV})|\mathcal{H}=0\right]$
$\leq$ $2\lambda_{n}+\frac{|\mathcal{U}||\mathcal{V}|}{4\mu_{n}^{2}n},$ (143)
and thus
$\tilde{P}_{UV}^{\otimes n}(\hat{\mathcal{E}}_{\bar{s}})$ $\geq$ $1-2\lambda_{n}-\frac{|\mathcal{U}||\mathcal{V}|}{4\mu_{n}^{2}n}.$ (144)
This probability can be decomposed into the contributions of the various type-classes. In fact, because $\hat{\mathcal{E}}_{\bar{s}}\subseteq\mathcal{T}_{\mu_{n}}^{(n)}(\tilde{P}_{UV})$ and because within a type-class all sequences are equally-likely:
$\hskip-14.22636pt\sum_{\begin{subarray}{c}\boldsymbol{\pi}_{UV}\colon\\ |\boldsymbol{\pi}_{UV}-\tilde{P}_{UV}|\leq\mu_{n}\end{subarray}}\tilde{P}_{UV}^{\otimes n}\left(\mathcal{T}_{0}^{(n)}(\boldsymbol{\pi}_{UV})\right)\frac{\left|(\hat{\mathcal{C}}_{\bar{s}}\times\hat{\mathcal{D}}_{\bar{s}})\cap\mathcal{T}_{0}^{(n)}(\boldsymbol{\pi}_{UV})\right|}{\left|\mathcal{T}_{0}^{(n)}(\boldsymbol{\pi}_{UV})\right|}$ (145)
$\geq$ $1-2\lambda_{n}-\frac{|\mathcal{U}||\mathcal{V}|}{4\mu_{n}^{2}n}.\hskip 113.81102pt$
Since the sum of probabilities
$\sum_{\begin{subarray}{c}\boldsymbol{\pi}_{UV}\colon\\ |\boldsymbol{\pi}_{UV}-\tilde{P}_{UV}|\leq\mu_{n}\end{subarray}}\tilde{P}_{UV}^{\otimes n}\left(\mathcal{T}_{0}^{(n)}(\boldsymbol{\pi}_{UV})\right)\leq 1,$ (146)
inequality (145) implies the existence of a type $\hat{\boldsymbol{\pi}}_{UV}$ within distance $\mu_{n}$ of $\tilde{P}_{UV}$ so that
$\frac{\left|(\hat{\mathcal{C}}_{\bar{s}}\times\hat{\mathcal{D}}_{\bar{s}})\cap\mathcal{T}_{0}^{(n)}(\hat{\boldsymbol{\pi}}_{UV})\right|}{\left|\mathcal{T}_{0}^{(n)}(\hat{\boldsymbol{\pi}}_{UV})\right|}\geq 1-2\lambda_{n}-\frac{|\mathcal{U}||\mathcal{V}|}{4\mu_{n}^{2}n}.$ (147)
We can use this bound on the cardinalities directly to obtain the desired lower bound on $\Delta_{\bar{s}}$. In fact, again using the fact that all sequences in a type-class have same probabilities under i.i.d. distributions:
$\Delta_{\bar{s}}$ (150)
$\geq$ $P_{UV}^{\otimes n}\left((\bar{\mathcal{C}}_{\bar{s}}\times\bar{\mathcal{D}}_{\bar{s}})\cap\mathcal{T}_{0}^{(n)}(\hat{\boldsymbol{\pi}}_{UV})\right)$
$=$ $\frac{\left|(\hat{\mathcal{C}}_{\bar{s}}\times\hat{\mathcal{D}}_{\bar{s}})\cap\mathcal{T}_{0}^{(n)}(\hat{\boldsymbol{\pi}}_{UV})\right|}{\left|\mathcal{T}_{0}^{(n)}(\boldsymbol{\pi}_{UV})\right|}\tilde{P}_{UV}^{\otimes n}\left(\mathcal{T}_{0}^{(n)}(\hat{\boldsymbol{\pi}}_{UV})\right)$
$\geq$ $\left(1-2\lambda_{n}-\frac{|\mathcal{U}||\mathcal{V}|}{4\mu_{n}^{2}n}\right)$
$\hskip 8.53581pt\cdot\ 2^{-nD(\hat{\boldsymbol{\pi}}_{UV}\|P_{UV})}\frac{1}{(n+1)^{|\mathcal{U}||\mathcal{V}|}}$
$\geq$ $\left(1-2\lambda_{n}-\frac{|\mathcal{U}||\mathcal{V}|}{4\mu_{n}^{2}n}\right)2^{-\max_{\hat{\boldsymbol{\pi}}_{UV}}nD(\hat{\boldsymbol{\pi}}_{UV}\|P_{UV})}$
$\ \cdot\ \frac{1}{(n+1)^{|\mathcal{U}||\mathcal{V}|}}$
where the maximization is over all types $\hat{\boldsymbol{\pi}}_{UV}$ that are of distance less than $\mu_{n}$ from $\tilde{P}_{UV}$.
This establishes the desired lower bound (138) and thus the proof of the lemma.

### v2

@@
 However, under $\mathcal{H}=0$, communication needs to be such that an external warden cannot detect the mere fact that communication is happening.
-Depending on the practical situation, it might be important that communication be covert under both hypothesie, or only in the case where there is no alert situation, i.e., under $\mathcal{H}=0$.
+Depending on the practical situation, it might be important that communication be covert under both hypotheses, or only in the case where there is no alert situation, i.e., under $\mathcal{H}=0$.
 In this paper, we consider the latter scenario.
+This asymmetric requirement is motivated by practical monitoring systems where no transmission should be detected under normal conditions $(\mathcal{H}=0)$, while when an abnormal event occurs $(\mathcal{H}=1)$, transmitting and thus being detected is acceptable.
 Implicitly, such a covertness constraint imposes that under $\mathcal{H}=0$, the sensor employs most of the time a dedicated zero-symbol that also represents the absence of communication.
@@
 where $n_{x^{n},y^{n}}(a,b)$ is the number of occurrences of the pair $(a,b)$ in $(x^{n},y^{n})$.
-We use $\mathcal{T}_{\mu}^{(n)}(P_{XY})$ to denote the jointly strongly typical set as defined in [21, Definition 2.9].
+We use $\mathcal{T}_{\mu}^{(n)}(P_{XY})$ to denote the jointly strongly typical set as defined in [7, Definition 2.9].
 II Problem Setup
@@
 $E_{1}$ $\triangleq$ $\min_{\begin{subarray}{c}\pi_{UV}\colon\\ \pi_{U}=P_{U}\\ \pi_{V}=P_{V}\end{subarray}}D(\pi_{UV}\|Q_{UV})$ (13a)
-$E_{2}$ $\triangleq$ $\min_{\begin{subarray}{c}\pi_{UV}:\ \\ \pi_{V}=P_{V},\;\pi_{U}\notin\bar{\mathcal{P}}_{U}(x_{1})\end{subarray}}D\!\big(\pi_{UV}\,\|\,Q_{UV}\big)$ (13b)
-$E_{3}$ $\triangleq$ $\max_{\begin{subarray}{c}x_{1}\in\mathcal{X}\setminus\{0\}\end{subarray}}\Big[\min_{\begin{subarray}{c}\pi_{UV}:\ \\ \pi_{V}=P_{V},\;\pi_{U}\in\bar{\mathcal{P}}_{U}\end{subarray}(x_{1})}D\!\big(\pi_{UV}\,\|\,Q_{UV}\big)$ (13c)
-$\hskip 85.35826pt+\ D(\Gamma_{Y|X=0}\|\Gamma_{Y|X=x_{1}})\Big],$
+$E_{2}(x_{1})$ $\triangleq$ $\min_{\begin{subarray}{c}\pi_{UV}:\ \\ \pi_{V}=P_{V},\;\pi_{U}\notin\bar{\mathcal{P}}_{U}(x_{1})\end{subarray}}D\big(\pi_{UV}\,\|\,Q_{UV}\big)$ (13b)
+$E_{3}(x_{1})$ $\triangleq$ $\min_{\begin{subarray}{c}\pi_{UV}:\ \\ \pi_{V}=P_{V},\;\pi_{U}\in\bar{\mathcal{P}}_{U}(x_{1})\end{subarray}}D\big(\pi_{UV}\,\|\,Q_{UV}\big)$ (13c)
+$\hskip 85.35826pt+\ D(\Gamma_{Y|X=0}\|\Gamma_{Y|X=x_{1}}),$
 […]
@@
     Otherwise
-    $\theta_{\textnormal{covert}}^{*}(\epsilon)$ $\geq$ $\min\big\{E_{2},E_{3}\big\}.$ (17)
+    $\theta_{\textnormal{covert}}^{*}(\epsilon)$ $\geq$ $\max_{\begin{subarray}{c}x_{1}\in\mathcal{X}\setminus\{0\}\end{subarray}}\min\big\{E_{2}(x_{1}),E_{3}(x_{1})\big\}.$ (17)
 Our achievability results do not require that the sensor and decision center use the shared secret key $S$.
@@
 $T_{U}(u):=\sum_{v}P_{V}(v)\,Q_{U|V}(u\mid v),\quad u\in\mathcal{U}.$ (18)
-We observe that if $T_{U}\notin\bar{\mathcal{P}}_{U}(x_{1})$ for all $x_{1}$, then $E_{2}=D(P_{V}\|Q_{V})$, which coincides with the local exponent at the decision center without communication from the sensor.
-In particular, $E_{2}=D(P_{V}\|Q_{V})$ when $T_{U}=P_{U}$ (as is the case for testing against independence where $Q_{UV}=P_{U}P_{V}$).
+We observe that if $T_{U}\notin\bar{\mathcal{P}}_{U}(x_{1})$ for all $x_{1}$, then $E_{2}(x_{1})=D(P_{V}\|Q_{V})$, which coincides with the local exponent at the decision center without communication from the sensor.
+In particular, $E_{2}(x_{1})=D(P_{V}\|Q_{V})$ when $T_{U}=P_{U}$ (as is the case for testing against independence where $Q_{UV}=P_{U}P_{V}$).
 For $T_{U}=P_{U}$ also $E_{1}=D(P_{V}\|Q_{V})$ and thus communication from the sensor cannot improve the Stein-exponent even for DMCs satisfying (15).
@@
 The second summand is strictly positive, except when $\pi_{U|V}(u|v)=Q_{U|V}(u|v)$ for all $u$ and all $v$ with $P_{V}(v)>0$.
-Thus, to deduce that $E_{1}>D(P_{V}\|Q_{V})$ or $E_{2}>D(P_{V}\|Q_{V})$ it suffices to prove that $\pi_{U|V}=Q_{U|V}$ is not a valid choice in the minimizations.
-Inspecting the two minimizations, we see that $\pi_{U|V}=Q_{U|V}$ is not a permissible choice in the minimization of $E_{1}$ when $T_{U}\neq P_{U}$ and of $E_{2}$ when $T_{U}\in\ \bar{\mathcal{P}}_{U}(x_{1})$ for some $x_{1}$.
-Notice also that $E_{3}>D(P_{V}\|Q_{V})$ by Assumptions (7). ∎
+Thus, to deduce that $E_{1}>D(P_{V}\|Q_{V})$ or $E_{2}(x_{1})>D(P_{V}\|Q_{V})$ it suffices to prove that $\pi_{U|V}=Q_{U|V}$ is not a valid choice in the minimizations.
+Inspecting the two minimizations, we see that $\pi_{U|V}=Q_{U|V}$ is not a permissible choice in the minimization of $E_{1}$ when $T_{U}\neq P_{U}$ and of $E_{2}(x_{1})$ when $T_{U}\in\ \bar{\mathcal{P}}_{U}(x_{1})$ for some $x_{1}$.
+Notice also that $E_{3}(x_{1})>D(P_{V}\|Q_{V})$ by Assumptions (7). ∎
 We now provide an example for which $\theta_{\textnormal{covert}}^{*}(\epsilon)>D(P_{V}\|Q_{V})$ even for fully-connected DMCs.
@@
 $E_{1}$ $=$ $D(P_{U}\|Q_{U})=0.7706$ (28)
-$E_{2}$ $=$ $\min_{\pi_{U}\notin\bar{\mathcal{P}}_{U}(1)}D(\pi_{U}\|Q_{U})=0.2095$ (29)
-$E_{3}$ $=$ $D(\Gamma_{Y|X}(\cdot|0)\|\Gamma_{Y|X}(\cdot|1))=0.1170.$ (30)
+$E_{2}(1)$ $=$ $\min_{\pi_{U}\notin\bar{\mathcal{P}}_{U}(1)}D(\pi_{U}\|Q_{U})=0.2095$ (29)
+$E_{3}(1)$ $=$ $D(\Gamma_{Y|X}(\cdot|0)\|\Gamma_{Y|X}(\cdot|1))=0.1170.$ (30)
 IV Proof of achievability results
@@
 $\cdot\ \Bigg(e^{-k\big(D(\pi\|\Gamma_{Z|X}(\cdot|x_{1}))-D(\pi\|\Gamma_{Z|X}(\cdot|0))\big)}-1\Bigg)^{2},$
-where in $(a)$ the second part of Lemma 1 and [21, Remark to Lemma 2.12].
+where in $(a)$ the second part of Lemma 1 and [7, Remark to Lemma 2.12].
 Since $k$ grows only sublinearly in $n$ and $\mu>0$, we can conclude that the right-hand side of (36) tends to 0 almost exponentially fast in $n$, thus satisfying the covertness constraint (11).
@@
 $\Pr\left[\hat{\mathcal{H}}=0,X^{n}=\bar{x}^{n}\Big|\mathcal{H}=0,S=\bar{s}\right]\geq\frac{1-\epsilon-\eta}{c|\tilde{\mathcal{X}}^{n}|}\bigg\}.$
-Notice that by Lemma 2 the cardinality of $|\tilde{\mathcal{X}}^{n}|$ grows subexponentially in $n$ and thus the ratio $\frac{1-\epsilon-\eta}{c|\tilde{\mathcal{X}}^{n}|}$ is of the form $e^{-n\phi_{n}}$ for a vanishing sequence $\phi_{n}$.
+Notice that by Lemma 2 the cardinality of $|\tilde{\mathcal{X}}^{n}|$ grows sub-exponentially in $n$ and thus the ratio $\frac{1-\epsilon-\eta}{c|\tilde{\mathcal{X}}^{n}|}$ is of the form $e^{-n\phi_{n}}$ for a vanishing sequence $\phi_{n}$.
 I.e., for any $\bar{s}\in\mathcal{\bar{S}}_{\bar{x}^{n}}$:
@@
 for
-$\phi_{n}:=\log\frac{c|\tilde{\mathcal{X}}^{n}|}{1-\epsilon-\eta}.$ (74)
+$\phi_{n}:=\frac{1}{n}\ln\frac{c|\tilde{\mathcal{X}}^{n}|}{1-\epsilon-\eta}.$ (74)
 Therefore, Lemma 3 implies that there exists a vanishing sequence $\eta_{n}$ so that for any $\bar{s}\in\mathcal{\bar{S}}_{\bar{x}^{n}}$:
@@
 $\stackrel{{\scriptstyle(a)}}{{\leq}}$ $\Pr\left[S\in\bar{\mathcal{S}}_{\bar{x}^{n}}\right]+\frac{1-\epsilon-\eta}{c|\tilde{\mathcal{X}}^{n}|}\bigg(1-\Pr\left[s\in\bar{\mathcal{S}}_{\bar{x}^{n}}\right]\bigg)$ (82)
-$=$ $\Pr\left[S\in\bar{\mathcal{S}}_{\bar{x}^{n}}\right]Bigg(1-\frac{1-\epsilon-\eta}{c|\tilde{\mathcal{X}}^{n}|}\Bigg)+\frac{1-\epsilon-\eta}{c|\tilde{\mathcal{X}}^{n}|},$ (83)
+$=$ $\Pr\left[S\in\bar{\mathcal{S}}_{\bar{x}^{n}}\right]\Bigg(1-\frac{1-\epsilon-\eta}{c|\tilde{\mathcal{X}}^{n}|}\Bigg)+\frac{1-\epsilon-\eta}{c|\tilde{\mathcal{X}}^{n}|},$ (83)
 where $(a)$ holds by the definition of $\bar{S}_{\bar{x}^{n}}$. Combining (71) and (83), we then obtain:
@@
 $\hat{\mathcal{D}}_{\bar{s}}$ $:=$ $\left\{\tilde{v}^{n}\colon\exists v^{n}\in\bar{\mathcal{D}}_{\bar{s}}\;\;{s.t.}\;\;d_{\text{H}}(\tilde{v}^{n},v^{n})\leq\ell_{n}\right\}.$ (120)
-By (118) and the blowing-up lemma [22, Remark on p. 446]:
+By (118) and the blowing-up lemma [12, Remark on p. 446]:
 $\tilde{P}_{U}^{\otimes n}(\hat{\mathcal{C}}_{\bar{s}})$ $\geq$ $1-\lambda_{n},$ (121a)
@@
 and the last inequality is obtained by simple counting arguments and because $2^{nH_{\textnormal{b}}(\ell_{n}/n)}$ upper bounds the set of all binary vectors with Hamming weight $\ell_{n}/n$.
-Let $\{\mu_{n}\}_{n=1}^{\infty}$ denote a sequence of positive numbers satisfying [^1] Condition (125b) ensures that the probability of the strongly typical set $\mathcal{T}_{\mu_{n}}^{(n)}(P_{XY})$ under $P_{XY}^{\otimes n}$ converges to one as $n\to\infty$; see [21, Remark following Lemma 2.12].
+Let $\{\mu_{n}\}_{n=1}^{\infty}$ denote a sequence of positive numbers satisfying [^1] Condition (125b) ensures that the probability of the strongly typical set $\mathcal{T}_{\mu_{n}}^{(n)}(P_{XY})$ under $P_{XY}^{\otimes n}$ converges to one as $n\to\infty$; see [7, Remark following Lemma 2.12].
 $\lim_{n\to\infty}\mu_{n}$ $=$ $0,$ (125a)